Orion Corporation has established the following standards for the prime costs of one unit of its chief product, dartboards. Standard Quantity Standard Price or Rate Standard Cost Direct material 8.00 pounds $ 1.30 per pound $ 10.40 Direct labor 0.50 hour $ 8.00 per hour 4.00 Total $ 14.40 During June, Orion purchased 285,000 pounds of direct material at a total cost of $407,550. The total wages for June were $156,060, 90 percent of which were for direct labor. Orion manufactured 34,000 dartboards during June, using 255,000 pounds of the direct material purchased in June and 18,360 direct-labor hours. Required: Compute the following variances for June. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "Favorable" or "Unfavorable". Select "None" and enter "0" for no effect (i.e., zero variance).)

1. Direct-material price variance
2. Direct-material quantity variance
3. Direct-material purchase price variance
4. Direct-labor rate variance
5. Direct-labor efficiency variance
Answer #1

1. Materials price variance = (Standard price - actual price) x Material used
= {1.30 - ($407,550 / 285000)} x 255000 = $33150 Unfavorable

2. Materials quantity variance = (Standard quantity - actual quantity used) x standard price
= {(34000*8) - 255000} x $1.30 = $22100 Favorable

3. Direct-material purchase price variance = (Standard price - actual price) x Material purchased
= {1.30 - ($407,550 / 285000)} x 285000 = $37050 Unfavorable

4. Labor rate variance = (Standard rate - actual rate) x Actual hours used
= {$8 - ($156,060*90% / 18360)} x 18360 = $6426 Favorable

Labor efficiency variance = (Standard hours - actual hours) x Standard rate
= {(34000*0.50) - 18360} x $8 = $10880 Unfavorable